For A = [-2 6
-3 4]
and B = [1 -4
-2 3]
Multiply AxB

Answer by ikleyn(52898)

.

            The process of filling matrix cells is shown below.

            Use your textbook as a guide and look at this post how it works.


Matrix A

        A1	A2
1	-2	6
2	-3	4

Matrix B

        B1	B2
1	1	-4
2	-2	3

c11 = -2 x 1 + 6 x (-2) = -14

         C1	C2
1	-14	0
2	0	0

c12 = -2 x (-4) + 6 x 3 = 26

         C1	C2
1	-14	26
2	0	0

c21 = -3 x 1 + 4 x (-2) = -11

         C1	C2
1	-14	26
2	-11	0

c22 = -3 x (-4) + 4 x 3 = 24

         C1	C2
1	-14	26
2	-11	24

Solved.